### v3

@@ 6.2. The nondeterminism hierarchy and GFMness
Let $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, $\mu\in{\mathcal{F}}({\mathcal{M}})$, be a finite memory strategy, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain where actions of ${\mathcal{M}}$ are resolved according to $\mu$. Notice that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $M\times S$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(M\times S)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(m,s)=\langle d,m^{\prime}\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\langle m,s\rangle,\langle m^{\prime},s^{\prime}\rangle)=0$. Let $r=\langle\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\langle m_{2},s_{2}\rangle,\ldots\rangle\in(M\times S)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$. Let $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{\omega}$ be a random run of ${\mathcal{A}}$ over the word $\tau(r)$ that is sampled by playing according to $\mu$ in ${\mathcal{M}}$. I.e., given $\langle m_{0},s_{0}\rangle,\langle m_{1},s_{1}\rangle,\ldots,\langle m_{k},s_{k}\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{A}}=\langle\langle m_{0},s_{0},q_{0}\rangle,\langle m_{1},s_{1},q_{1}\rangle,\ldots\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{A}}$ with a set of positions $M\times S\times Q_{\mathcal{A}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$ can be expressed as follows:
[…]
@@
Since ${\mathcal{C}}_{\mathcal{A}}$ is a finite Markov chain, it holds that with probability 1 the random process $r\oplus r_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component and traverses all its states. We say that an ergodic component $C\subseteq M\times S\times Q_{\mathcal{A}}$ of ${\mathcal{C}}_{\mathcal{A}}$ is $\alpha_{\mathcal{A}}$-accepting if the projection of $C$ onto $Q_{\mathcal{A}}$ intersects $\alpha_{\mathcal{A}}$. I.e., $C$ is $\alpha_{\mathcal{A}}$-accepting if there exist $s\in S$, $m\in M$ and $q\in\alpha_{\mathcal{A}}$ such that $\langle m,s,q\rangle\in C$. Thus, if $r\oplus r_{\mathcal{A}}$ eventually enters an $\alpha_{\mathcal{A}}$-accepting ergodic component $C$, then with probability 1 the process traverses all the positions in $C$ infinitely often, and in particular with probability 1 $r_{\mathcal{A}}$ is an accepting run of ${\mathcal{A}}$ over the generated word $\tau(r)$. Thus, since with probability 1 the process of ${\mathcal{C}}_{\mathcal{A}}$ eventually enters an ergodic component, it follows that the probability that $\tau(r)\in L({\mathcal{A}})$ but $r_{\mathcal{A}}$ is rejecting equals the probability that $r\oplus r_{\mathcal{A}}$ enters an ergodic component that is not $\alpha_{\mathcal{A}}$-accepting and $\tau(r)\in L({\mathcal{A}})$. We prove that if $r\oplus r_{\mathcal{A}}$ enters an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting then $\tau(r)\notin L({\mathcal{A}})$. Note that we claim that the implication $C$ not $\alpha_{\mathcal{A}}$-accepting implies that $\tau(r)\notin L({\mathcal{A}})$, is surely, and not only almost-surely. Consider an ergodic component $C$ that is not $\alpha_{\mathcal{A}}$-accepting, and assume that $r\oplus r_{\mathcal{A}}$ is such that $\langle m_{k},s_{k},q_{k}\rangle\in C$ for some $k\geq 0$. Assume by way of contradiction that $\tau(r)\in L({\mathcal{A}})$. Since ${\mathcal{A}}$ is SD, it is possible to extend the finite run $q_{0},q_{1},\ldots q_{k}$ of ${\mathcal{A}}$ over $\tau(s_{1}),\tau(s_{2}),\ldots,\tau(s_{k})$ into an accepting run of ${\mathcal{A}}$ over $\tau(r)$. In particular there exist $j>k$ and $q^{\prime}_{k+1},q^{\prime}_{k+2},\ldots,q^{\prime}_{j}\in(Q_{\mathcal{A}})^{*}$, such that $q^{\prime}_{j}\in\alpha_{\mathcal{A}}$ and $q_{0},\ldots,q_{k},q^{\prime}_{k+1},\ldots,q^{\prime}_{j}$ is a finite run of ${\mathcal{A}}$ over $\tau(s_{1}),\ldots,\tau(s_{j})$. By definition of the stochastic transition function of ${\mathcal{C}}_{\mathcal{A}}$, it follows that there is a positive probability to move to $\langle m_{k+1},s_{k+1},q^{\prime}_{k+1}\rangle$ from $\langle m_{k},s_{k},q_{k}\rangle$. Similarly, it follows by induction that for all $k+1<i\leq j$ there is a positive probability to move to $\langle m_{i},s_{i},q^{\prime}_{i}\rangle$ from $\langle m_{i-1},s_{i-1},q^{\prime}_{i-1}\rangle$. In particular $\langle m_{j},s_{j},q^{\prime}_{j}\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{A}}$-accepting. Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{A}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{A}}$ over $\tau(r)$. It is not hard to see that ${\mathcal{C}}_{\mathcal{A}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{A}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$. We proved that there is a zero probability that $\tau(r)\in L({\mathcal{A}})$ and $r_{\mathcal{A}}$ is rejecting. Thus, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$.
We can now conclude with the desired connection between SDness and GFMness.
@@
We first examine whether SDness implies GFMness, and show that the answer depends on the acceptance condition.
@@
The co-Büchi case follows from Theorems 11 and 15. For Büchi automata, consider an SD-NBW ${\mathcal{A}}=\langle\Sigma,Q_{\mathcal{A}},q_{0},\delta_{\mathcal{A}},\alpha_{\mathcal{A}}\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$. Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ given by [11]. Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$. Since $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ holds for all ${\mathcal{A}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$.
@@ Appendix A Determinization of SD-NBWs
Consider states $q\in Q$ and $S\in Q^{\prime}$, a letter $\sigma\in\Sigma$, and transitions $\langle q,\sigma,q^{\prime}\rangle$ and $\langle S,\sigma,S^{\prime}\rangle$ of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, respectively. If $q$ is ${\mathcal{A}}$-equivalent to the states in $S$, then $q^{\prime}$ is ${\mathcal{A}}$-equivalent to the states in $S^{\prime}$.
@@
Consider a state $S$ of ${\mathcal{A}}^{\prime}$ and a letter $\sigma\in\Sigma$. If $\langle S,\sigma,S^{\prime}\rangle\in\Delta^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, then all the $\sigma$-successors of a state $s\in S$ are in $S^{\prime}\setminus\alpha$.
@@
Consider the DAG $G=\langle V,E\rangle$, where
@@
  • $E\subseteq\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})\times(S_{j_{i+1}}\times\{i+1\})$ is such that for all $i\geq 0$, it holds that $E(\langle s^{\prime},i\rangle,\langle s,i+1\rangle)$ iff there is a finite run from $s^{\prime}$ to $s$ over $w[j_{i}+1,j_{i+1}]$. Then, we label this edge by the run from $s^{\prime}$ to $s$.
By the definition of ${\mathcal{A}}^{\prime}$, for every $j\geq 0$ and state $s_{j+1}\in S_{j+1}$, there is a state $s_{j}\in S_{j}$ such that $\langle s_{j},\sigma_{j},s_{j+1}\rangle\in\Delta$. Thus, it follows by induction that for every $i\geq 0$ and state $s_{i+1}\in S_{j_{i+1}}$, there is a state $s_{i}\in S_{j_{i}}$ such that there is a finite run from $s_{i}$ to $s_{i+1}$ on $w[j_{i}+1,j_{i+1}]$. Thus, the DAG $G$ has infinitely many reachable vertices from the vertex $\langle q_{0},0\rangle$. Also, as the nondeterminism degree of ${\mathcal{A}}$ is finite, so is the branching degree of $G$. Thus, by König’s Lemma, $G$ includes an infinite path, and the labels along the edges of this path define a run of ${\mathcal{A}}$ on $w$. Since for all $i\geq 1$, the state $S_{j_{i}}$ is in $\alpha^{\prime}$, and so all the states in $S_{j_{i}}$ are in $\alpha$, this run is accepting, and we are done.
For the other direction, assume that $w=\sigma_{1}\cdot\sigma_{2}\cdots\in L({\mathcal{A}})$, and let $r=r_{0},r_{1},\ldots$ be an accepting run of ${\mathcal{A}}$ on $w$. Let $S_{0},S_{1},S_{2}\ldots$ be the run of ${\mathcal{A}}^{\prime}$ on $w$, and assume, by way of contradiction, that there is a position $j\geq 0$ such that $S_{j},S_{j+1},\ldots$ is an $\alpha$-free run on the suffix $w[j+1,\infty]$. Then, an iterative application of Proposition 20 implies that all the runs of a state $s_{j}\in S_{j}$ on $w[j+1,\infty]$ are $\alpha$-free in ${\mathcal{A}}$. Also, an iterative application of Proposition 19 implies that $r_{j}\sim_{\mathcal{A}}s_{j}$, and since $r$ is an accepting run of ${\mathcal{A}}$, it holds that ${\mathcal{A}}^{s_{j}}$ has an accepting run on $w[j+1,\infty]$, and we have reached a contradiction.
@@
Assume by way of contradiction that there are reachable states $S\in\alpha^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, and an infinite run $r_{{\mathcal{A}}^{\prime}}=S_{0},S_{1},S_{2},\ldots$ that visits both $S$ and $S^{\prime}$ infinitely often. Recall that a reachable state in $Q^{\prime}$ contains only $\alpha$-states of ${\mathcal{A}}$ or only $\bar{\alpha}$-states of ${\mathcal{A}}$. Hence, $S^{\prime}$ contains only $\bar{\alpha}$-states of ${\mathcal{A}}$. As in the proof of Proposition 21, the run $r_{{\mathcal{A}}^{\prime}}$ induces an infinite run $r_{\mathcal{A}}=s_{0},s_{1},s_{2},\ldots$, where for all positions $j\geq 0$, it holds that $s_{j}\in S_{j}$. Since the run $r_{{\mathcal{A}}^{\prime}}$ visits $S$ infinitely often, then $r_{\mathcal{A}}$ visits infinitely many $\alpha$-states. Likewise, since $r_{{\mathcal{A}}^{\prime}}$ visits $S^{\prime}$ infinitely often, then $r_{\mathcal{A}}$ also visits infinitely many $\bar{\alpha}$-states. This contradicts the weakness of ${\mathcal{A}}$, and we are done.

### v4

@@ 6.2. The nondeterminism hierarchy and GFMness
Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in ${\mathcal{M}}$ according to $\mu$. Note that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise. Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$. Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of ${\mathcal{N}}$ over the word $\tau(r)$ that is sampled when playing in ${\mathcal{M}}$ with the startegy $\mu$. I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
$\mathrm{Pr}_{{\mathcal{C}}_{\mathcal{N}}}(\left\langle s,m,q\right\rangle,\left\langle s^{\prime},m^{\prime},q^{\prime}\right\rangle)=\begin{cases}|\delta_{\mathcal{N}}(q,\tau(s^{\prime}))|^{-1}\cdot\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)&:q^{\prime}\in\delta_{\mathcal{N}}(q,\tau(s^{\prime}))\\ 0&:\mbox{otherwise}\end{cases}$
@@
Since ${\mathcal{C}}_{\mathcal{N}}$ is a finite Markov chain, it holds that with probability $1$, the random process $r\oplus r_{\mathcal{N}}$ of ${\mathcal{C}}_{\mathcal{N}}$ eventually enters an ergodic component and traverses all its states. We say that an ergodic component $C\subseteq S\times M\times Q_{\mathcal{N}}$ of ${\mathcal{C}}_{\mathcal{N}}$ is $\alpha_{\mathcal{N}}$-accepting if the projection of $C$ onto $Q_{\mathcal{N}}$ intersects $\alpha_{\mathcal{N}}$. I.e., $C$ is $\alpha_{\mathcal{N}}$-accepting if there exist $s\in S$, $m\in M$ and $q\in\alpha_{\mathcal{N}}$ such that $\left\langle s,m,q\right\rangle\in C$. Thus, if $r\oplus r_{\mathcal{N}}$ eventually enters an $\alpha_{\mathcal{N}}$-accepting ergodic component $C$, then with probability 1 the process traverses all the positions in $C$ infinitely often, and in particular with probability 1 $r_{\mathcal{N}}$ is an accepting run of ${\mathcal{N}}$ over the generated word $\tau(r)$. Since with probability 1 the process of ${\mathcal{C}}_{\mathcal{N}}$ eventually enters an ergodic component, it follows that the probability that $\tau(r)\in L({\mathcal{N}})$ but $r_{\mathcal{N}}$ is rejecting equals the probability that $r\oplus r_{\mathcal{N}}$ enters an ergodic component that is not $\alpha_{\mathcal{N}}$-accepting and $\tau(r)\in L({\mathcal{N}})$. We prove that if $r\oplus r_{\mathcal{N}}$ enters an ergodic component $C$ that is not $\alpha_{\mathcal{N}}$-accepting then $\tau(r)\notin L({\mathcal{N}})$. Note that we claim that the implication $C$ not $\alpha_{\mathcal{N}}$-accepting implies that $\tau(r)\notin L({\mathcal{N}})$, is surely, and not only almost-surely. Consider an ergodic component $C$ that is not $\alpha_{\mathcal{N}}$-accepting, and assume that $r\oplus r_{\mathcal{N}}$ is such that $\left\langle s_{k},m_{k},q_{k}\right\rangle\in C$ for some $k\geq 0$. Assume by way of contradiction that $\tau(r)\in L({\mathcal{N}})$. Since ${\mathcal{N}}$ is SD, it is possible to extend the finite run $q_{0},q_{1},\ldots q_{k}$ of ${\mathcal{N}}$ over $\tau(s_{1}),\tau(s_{2}),\ldots,\tau(s_{k})$ into an accepting run of ${\mathcal{N}}$ over $\tau(r)$. In particular there exist $j>k$ and $q^{\prime}_{k+1},q^{\prime}_{k+2},\ldots,q^{\prime}_{j}\in(Q_{\mathcal{N}})^{*}$, such that $q^{\prime}_{j}\in\alpha_{\mathcal{N}}$ and $q_{0},\ldots,q_{k},q^{\prime}_{k+1},\ldots,q^{\prime}_{j}$ is a finite run of ${\mathcal{N}}$ over $\tau(s_{1}),\ldots,\tau(s_{j})$. By definition of the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$, it follows that there is a positive probability to move to $\left\langle s_{k+1},m_{k+1},q^{\prime}_{k+1}\right\rangle$ from $\left\langle s_{k},m_{k},q_{k}\right\rangle$. Similarly, it follows by induction that for all $k+1<i\leq j$ there is a positive probability to move to $\left\langle s_{i},m_{i},q^{\prime}_{i}\right\rangle$ from $\left\langle s_{i-1},m_{i-1},q^{\prime}_{i-1}\right\rangle$. In particular $\left\langle s_{j},m_{j},q^{\prime}_{j}\right\rangle\in C$, in contradiction to the assumption that $C$ is not $\alpha_{\mathcal{N}}$-accepting.
Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{N}}$ over $\tau(r)$. It is not hard to see that ${\mathcal{C}}_{\mathcal{N}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{N}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$. We proved that there is a zero probability that $\tau(r)\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting. Thus, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$.
@@
We can now conclude with the desired connection between SDness and GFMness. We first examine whether SDness implies GFMness, and show that the answer depends on the acceptance condition.
@@
The co-Büchi case follows from Theorems 11 and 15. For Büchi automata, consider an SD-NBW ${\mathcal{N}}=\left\langle\Sigma,Q_{\mathcal{N}},q_{0},\delta_{\mathcal{N}},\alpha_{\mathcal{N}}\right\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$. Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ given by [CJH03]. Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$. Since $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ holds for all ${\mathcal{N}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
@@
We show that ${\mathcal{N}}$ is GFM. Consider an MDP ${\mathcal{M}}$, and let $\mu:S\times M\to dist(A)\times M$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$. Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
To conclude the GFMness of ${\mathcal{N}}$, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$. Recall that a strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in ${\mathcal{M}}$ and in ${\mathcal{N}}$ simultaneously. In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in ${\mathcal{M}}$, while in the latter it should choose a $\tau(s)$-succsessor of the current state $q$ in ${\mathcal{N}}$. The strategy $\mu^{\prime}$ resolves the first according to to $\mu$, and hence ignores the ${\mathcal{N}}$ component. Then, for the ${\mathcal{N}}$ component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of ${\mathcal{N}}$ on $w_{\mathcal{C}}$ as follows. The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in ${\mathcal{C}}$; in particular, if the ergodic SCC that is reached in ${\mathcal{C}}$ contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$. In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
As the MC ${\mathcal{C}}$ is finite, then with probability 1 an ergodic component of ${\mathcal{C}}$ is reached and all of its states are traversed infinitely often. Hence, with probability 1, we have one of the following outcomes. Either the ergodic component that is reached does not contains an $a$-labeled state, or it contains an $a$-labeled state. In the former case, the run $r_{\mathcal{N}}$ is accepting. Indeed, according to $\mu^{\prime}$, the run $r_{\mathcal{N}}$ moves to $q_{acc}$ once the ergodic component that has no $a$-labeled states is reached, and thus $r_{\mathcal{N}}$ eventually gets stuck at $q_{acc}$ while reading a suffix of $w_{\mathcal{C}}$ that has no $a$’s. In the latter case, $r_{\mathcal{N}}$ is stuck in $q_{0}$ and is thus rejecting, and with probability 1 $w_{\mathcal{C}}$ has infinitely many $a$’s. Hence, in the latter case there is a zero probability that $w_{\mathcal{C}}\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting, and we are done.
@@ Appendix A Determinization of SD-NBWs
Consider states $q\in Q$ and $S\in Q^{\prime}$, a letter $\sigma\in\Sigma$, and transitions $\left\langle q,\sigma,q^{\prime}\right\rangle$ and $\left\langle S,\sigma,S^{\prime}\right\rangle$ of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, respectively. If $q$ is ${\mathcal{A}}$-equivalent to the states in $S$, then $q^{\prime}$ is ${\mathcal{A}}$-equivalent to the states in $S^{\prime}$.
@@
Consider a state $S$ of ${\mathcal{A}}^{\prime}$ and a letter $\sigma\in\Sigma$. If $\left\langle S,\sigma,S^{\prime}\right\rangle\in\Delta^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, then all the $\sigma$-successors of a state $s\in S$ are in $S^{\prime}\setminus\alpha$.
@@
Consider the DAG $G=\left\langle V,E\right\rangle$, where
@@
  • $E\subseteq\bigcup_{i\geq 0}(S_{j_{i}}\times\{i\})\times(S_{j_{i+1}}\times\{i+1\})$ is such that for all $i\geq 0$, it holds that $E(\left\langle s^{\prime},i\right\rangle,\left\langle s,i+1\right\rangle)$ iff there is a finite run from $s^{\prime}$ to $s$ over $w[j_{i}+1,j_{i+1}]$. Then, we label this edge by the run from $s^{\prime}$ to $s$.
By the definition of ${\mathcal{A}}^{\prime}$, for every $j\geq 0$ and state $s_{j+1}\in S_{j+1}$, there is a state $s_{j}\in S_{j}$ such that $\left\langle s_{j},\sigma_{j},s_{j+1}\right\rangle\in\Delta$. Thus, it follows by induction that for every $i\geq 0$ and state $s_{i+1}\in S_{j_{i+1}}$, there is a state $s_{i}\in S_{j_{i}}$ such that there is a finite run from $s_{i}$ to $s_{i+1}$ on $w[j_{i}+1,j_{i+1}]$. Thus, the DAG $G$ has infinitely many reachable vertices from the vertex $\left\langle q_{0},0\right\rangle$. Also, as the nondeterminism degree of ${\mathcal{A}}$ is finite, so is the branching degree of $G$. Thus, by König’s Lemma, $G$ includes an infinite path, and the labels along the edges of this path define a run of ${\mathcal{A}}$ on $w$. Since for all $i\geq 1$, the state $S_{j_{i}}$ is in $\alpha^{\prime}$, and so all the states in $S_{j_{i}}$ are in $\alpha$, this run is accepting, and we are done.
For the other direction, assume that $w=\sigma_{1}\cdot\sigma_{2}\cdots\in L({\mathcal{A}})$, and let $r=r_{0},r_{1},\ldots$ be an accepting run of ${\mathcal{A}}$ on $w$.
Let $S_{0},S_{1},S_{2}\ldots$ be the run of ${\mathcal{A}}^{\prime}$ on $w$, and assume, by way of contradiction, that there is a position $j\geq 0$ such that $S_{j},S_{j+1},\ldots$ is an $\alpha$-free run on the suffix $w[j+1,\infty]$. Then, an iterative application of Proposition 20 implies that all the runs of a state $s_{j}\in S_{j}$ on $w[j+1,\infty]$ are $\alpha$-free in ${\mathcal{A}}$. Also, an iterative application of Proposition 19 implies that $r_{j}\sim_{\mathcal{A}}s_{j}$, and since $r$ is an accepting run of ${\mathcal{A}}$, it holds that ${\mathcal{A}}^{s_{j}}$ has an accepting run on $w[j+1,\infty]$, and we have reached a contradiction.
@@
Assume by way of contradiction that there are reachable states $S\in\alpha^{\prime}$ and $S^{\prime}\notin\alpha^{\prime}$, and an infinite run $r_{{\mathcal{A}}^{\prime}}=S_{0},S_{1},S_{2},\ldots$ that visits both $S$ and $S^{\prime}$ infinitely often. Recall that a reachable state in $Q^{\prime}$ contains only $\alpha$-states of ${\mathcal{A}}$ or only $\bar{\alpha}$-states of ${\mathcal{A}}$. Hence, $S^{\prime}$ contains only $\bar{\alpha}$-states of ${\mathcal{A}}$.
As in the proof of Proposition 21, the run $r_{{\mathcal{A}}^{\prime}}$ induces an infinite run $r_{\mathcal{A}}=s_{0},s_{1},s_{2},\ldots$, where for all positions $j\geq 0$, it holds that $s_{j}\in S_{j}$. Since the run $r_{{\mathcal{A}}^{\prime}}$ visits $S$ infinitely often, then $r_{\mathcal{A}}$ visits infinitely many $\alpha$-states. Likewise, since $r_{{\mathcal{A}}^{\prime}}$ visits $S^{\prime}$ infinitely often, then $r_{\mathcal{A}}$ also visits infinitely many $\bar{\alpha}$-states. This contradicts the weakness of ${\mathcal{A}}$, and we are done.